OpenStudy (chrk_1):

In section 90, the professor explains that X^X = e ^ xlnx. Could some one explains little details about this and how is it possible to write in this way

OpenStudy (phi):

by definition, ln x = some exponent (call it "a") such that e^a = x that is, given ln x = a e^a = x or using a = ln x \[ e^{\ln x} = x\] this is an identity

OpenStudy (phi):

if we write \[ x = e^{\ln x} \] then \[ x^x = \left( e^{\ln x} \right)^x \] there is a property of exponents: \[ \left(a^b\right)^c = a^{b c}\] which allows us to write \[ x^x = \left( e^{\ln x} \right)^x = e^{x\ln x} \]